# Quantum Bit Threads and the Entropohedron

### v1

@@ 2.6 Changing the cutoff
Now let us consider what happens if we change the cutoff $\epsilon$. Recall that, as usual in physics, the cutoff is not a physical parameter but a calculational device we impose. Under this change, the area term $|\eth r|/\epsilon^{D-2}$ and therefore the bulk entropy change, while $G_{\rm N}$ makes a compensating change so that the physical quantity $S_{\rm gen}(r)$ is invariant Susskind:1994sm ; Cooperman:2013iqr . (Recall that $G_{\rm N}$ is the bare parameter appearing in the Lagrangian.) In particular, if we decrease $\epsilon$, then both $S_{\rm b}$ and $G_{\rm N}$ increase, so the divergence constraint becomes looser while the norm bound becomes tighter. The flux of $v$ through $A$ does not change, but what happens to that flux in the bulk changes, as less of it can go through the QES and more of it must jump across.
@@ 2.7 Theorems & proofs
In this subsection, we state and prove several theorems that were used in the rest of the section. These proofs use the same tools of convex analysis, especially convex relaxation and dualization, as in the Riemannian max flow-min cut theorem and related theorems, which may be found in Headrick:2017ucz . Therefore, we will be brief. We will also not make any attempt to be rigorous, particularly as regards the measure theory and functional analysis aspects of the proofs.
@@ 2.7.2 Theorems
The proof is similar to that in Rolph:2021hgz for the corresponding theorem for loose flows. We write the RHS of (67) as a concave program:
@@
This theorem was proved in Rolph:2021hgz . Here, we give an alternate proof that involves dualizing from surfaces to flows, similar to the second proof of theorem 2.4. On a technical note, this proof requires an additional assumption, that there is a large volume near $\partial\Sigma$, such as when $\Sigma$ is the region within a large radial cutoff surface of an asymptotically hyperbolic manifold.
@@ 2.8 Multiflows
From the definition, each component vector field $v_{ij}$ is itself a classical flow. More than that, any linear combination $\sum_{i<j}c_{ij}v_{ij}$ where the coefficients obey $|c_{ij}|\leq 1$ is a classical flow. In particular, for any individual boundary region $A_{i}$, we can define the flow $v_{i}:=\sum_{j}v_{ij}$; its flux $\int_{A_{i}}v_{i}$ is bounded above by $S(A_{i})$. The following theorem, proved in Cui:2018dyq , shows that there exists a classical multiflow that simultaneously saturates these bounds for all $i$:
@@ 3.2.2 The strict constraint is saturated on the entanglement wedge
We can actually show something more general that implies what we want: if $\int_{A}v=S_{\rm gen}({\eth r})$, for any $\eth r$} homologous to $A$, then (112) is saturated for that $\eth r$}, and $v$ is perpendicular to $\eth r$}. Applying (112) to any $\eth r$} homologous to $A$ and with $\int_{A}v=S_{\rm gen}({\eth r})$ gives
$\int_{\eth r}|v|+\left|\int_{\eth r}v-S_{\rm gen}({\eth r})\right|\leq S_{\rm gen}({\eth r}).$ (117)
@@
$\int_{\eth r}|v|+\int_{\eth r}v\leq 2S_{\rm gen}({\eth r}),$ (119)
and if either one of these is saturated then (117) is saturated. Only (118) can be saturated, and it forces $v$ on $\eth r$} to be parallel to the normal.
@@ 3.2.3 Constraints implied by the strict constraint
$\begin{split}|v(x)|&\leq\inf_{\eth r\ni x}\frac{S_{\rm gen}(\eth r)}{|\eth r|}\\ &=\frac{1}{4G_{N}}+DS_{b}(x),\qquad DS_{b}(x)=\inf_{\eth r\ni x}\frac{S_{b}(\eth r)}{|\eth r|}\end{split}$ (124)
@@ 3.3 Divergenceless, loose, cutoff-independent flows
The divergenceless, loose, cutoff-independent prescription is (110). These divergenceless flows have similarities and differences with respect to the strict flows. The divergenceless flow lines cannot start and end at points in the bulk. From their definition, the flows obey (122), just like the strict flows, but only for $r\in\mathcal{R}_{A}$. The divergenceless flows do obey the same local norm bound (124) as the strict flow, but we will have to derive it another way, because in this prescription we cannot apply (122) to arbitrary bulk regions.
@@ 3.4.1 Loose flow prescription
$\sup_{v}\int_{A}v\qquad\text{subject to }\int_{{\eth r}}|v|-\int_{r}\nabla\cdot v\leq S_{\rm gen}({\eth r})$ (128)
@@
$L[v,\mu]=\int_{A}v-\int_{\partial\Sigma}\psi v-\int_{\mathcal{R}_{A}}d\mu\left(\int_{\Sigma}\left(d\chi_{r}\cdot v+|v||d\chi_{r}|\right)+S_{\rm gen}(r)\right),\qquad\psi:=\int_{\mathcal{R}_{A}}d\mu\chi_{r}.$ (130)
@@ 3.4.3 Strict flow prescription: surface to flow dualisation
Here we will dualise from the QES prescription to the strict flow prescription. Since we already proved the equivalence between the prescriptions in section 3.4.2, this derivation is redundant, but we include it to show how it can be done because surface-to-flow dualisations are needed when deriving covariant prescriptions CQBT .
@@ 3.4.4 Divergenceless loose flow prescription
To understand when (147) is QES-equivalent, it is useful to note that an implication of its constraints is
@@
Now, if $A$ has a boundary, then (147) is equivalent to the QES prescription: from (148), if $|d\chi_{A}|=1$ at any point, then we must have $\int_{\mathcal{R}_{A}}d\mu\geq 1$, and so (147) is lower bounded by $S(A)$. But it is also upper bounded by $S(A)$, because $\lambda=\chi_{r_{QES}}$ and $\mu(r)=\delta(r-r_{QES})$ is always feasible. So, (147) gives $S(A)$, if $A$ has a boundary.
@@ 4 Quantum thread distributions
Thread distributions, introduced in Headrick:2022nbe , are an alternative mathematical representation of bit threads, replacing the flow $v$ with a measure $\upmu$ over bulk curves connecting boundary points. Although they were defined both in the static and covariant settings, here we will focus (as we do throughout this paper) with the static case, and explain a possible method to reproduce the QES formula using thread distributions. Unfortunately, we will have to leave several key statements as conjectures.
@@ 4.1 Entanglement pair functions
In upgrading classical TDs to quantum ones, we will need the concept of an entanglement pair function, which represents the bulk entropies and plays the same role for quantum TDs that the entanglement density function (EDF), defined and studied in subsection 2.3 and section 5, play for quantum flows. We define an entanglement pair function (EPF) as a non-negative function $g$ on $\Sigma^{2}:=\Sigma\times\Sigma$ such that [^11] Because our application of EPFs will be in holography, in defining EPFs, we are using the notation corresponding to bulk entropies in holography. However, the concept can be applied to any state of a field theory for which entropies of spatial regions are defined, or even more generally, to any multiparty quantum state. Note also that (151) is a property shared by the so-called two-point partial entanglement entropy $\mathcal{I}(x,y)$, defined in Lin:2023rxc .
@@ 4.4 Double holography
In the double holography setting, making use of the above mapping from classical TDs to EPFs, we can also prove conjecture 4.1. Namely, starting from an EDF $f$ on $\Sigma$, we lift it to a classical flow $\tilde{v}$ in $\tilde{\Sigma}$ using theorem 2.2; we then map $\tilde{v}$ to a classical TD $\tilde{\upmu}$ on $\tilde{\Sigma}$ via the flow lines of $\tilde{v}$, as described in subsection 5.1.1 of Headrick:2022nbe (except that, rather than restricting to threads connecting specified boundary regions, we keep all threads connecting any two boundary points, and we retain their orientations); finally, using (172) we map $\tilde{\upmu}$ to an EPF $g$. Through this series of transformations, $f(x)$ becomes the boundary flux $\tilde{n}\cdot\tilde{v}(x)$, which in turn becomes the number of threads beginning (if $f(x)>0$) or ending (if $f(x)<0$) at $x$, finally implying that $g$ obeys $\int_{\Sigma}dy\,g(x,y)=|f(x)|$.
@@ 5 The entropohedron
In the definition of strict quantum flows given in section 2, the concept of entanglement distribution function (EDF) played a key role. In that setting, the relevant entropies were those of spatial regions of the bulk in a holographic spacetime. The notion of EDF is also applicable for a quantum state on a finite number of parties. For $N$ parties, the EDFs elegantly repackage the entropy function, which is a function of $2^{N}-1$ variables (or equivalently the entropy vector, a vector in $2^{N}-1$ dimensions), into a convex polytope in just $N$ dimensions, which we will call the entropohedron. Whereas the entropy function treats all sets of parties on an equal footing, the entropohedron makes use of the local structure, i.e. which sets of parties are contained within which other sets. It also builds in important properties of entropies such as strong subadditivity. In this section, we give a self-contained presentation of this object and its properties. In this finite-dimensional setting, we will be able to prove the relevant properties rigorously, as no analysis is required.
The entropohedron is closely related to the symmetric submodular polytope studied in the theory of submodular functions, and many of the properties and relations we will discuss have also been proven for that object (see for example bach2013learning ). However, there is a difference in that the monotonicity property $S(AB)\geq S(A)$ usually imposed in that setting is replaced here by the weak monotonicity property (180). The simple relation to flows on graphs, discussed in subsection 5.5, has also not been noted before as far we know.
@@
5.1 Definitions & basic properties
Let X{X} be a finite set and $2^{X}$ its power set (the set of all subsets of X{X}). We will use $x,y,\ldots$ to represent elements of X{X} and $A,B,\ldots$ to represent elements of $2^{X}$. We denote by $\mathbf{R}^{X}$ the |X||{X}|-dimensional real vector space with basis vectors labelled by elements of X{X}.
@@ 5.1 Definitions & basic properties
An entropy function [^15] SS is also sometimes called an entropy vector, as it may be considered a vector in $\mathbf{R}^{2^{X}\setminus\{\emptyset\}}$, where we remove the empty set from the basis since we require $S(\emptyset)=0$., [^16] The strong subadditivity condition (179) is also called submodularity. In the literature on submodular functions, in particular in defining the symmetric submodular polytope, the condition (178) is also usually imposed, along with monotonicity, $S(AB)\geq S(A)$ bach2013learning . Monotonicity is obeyed by the entropies of classical probability distributions; however, for quantum states it is replaced by the weak monotonicity condition (180). for X{X} is a function $S:2^{X}\to[0,\infty)$, satisfying the following conditions:
@@
$\forall\,A,B\in 2^{X}\,,\quad S(A)+S(B)$ $\geq S(A\cap B)+S(A\cup B)\qquad\text{(SSA)}$ (179)
$\forall\,A,B\in 2^{X}\,,\quad S(A)+S(B)$ $\geq S(A\setminus B)+S(B\setminus A)\qquad\text{(WM)}\,.$ (180)
The entropies of a multipartite quantum system in a fixed state obey (178)–(179), hence the name “entropy function”. [^17] The name “entropy function” is a bit of an abuse, since not every entropy function represents the entropies of a quantum state. In particular, it is known that the entropies of quantum states on at least four parties obey a further set of constrained inequalities Linden:2004ebt ; Cadney:2011vix . For this reason, we will sometimes use the language of “states” and “parties”. The only aspects of the state that are relevant for our purposes, however, are the entropies. Another natural source of functions obeying these properties is minimal cuts on weighted graphs. We will discuss these further in subsection 5.5. Of course, minimal cuts on graphs are related to quantum states via the RT formula.
@@
Given an entropy function $S$, an entanglement distribution function (EDF) is a function $f:X\to\mathbf{R}$ such that, for all $A\in 2^{X}$,
@@
The entropohedron $F_{S}$ is the set of EDFs of $S$.
The following properties of the entropohedron follow directly from the definition:

### v2

@@ 2.6 Changing the cutoff
Now let us consider what happens if we change the cutoff $\epsilon$. Recall that, as usual in physics, the cutoff is not a physical parameter but a calculational device we impose. Under this change, the area term $|\eth r|/\epsilon^{D-2}$ and therefore the bulk entropy change, while $G_{\rm N}$ makes a compensating change so that the physical quantity $S_{\rm gen}(r)$ is invariant Susskind and Uglum (1994); Cooperman and Luty (2014). (Recall that $G_{\rm N}$ is the bare parameter appearing in the Lagrangian.) In particular, if we decrease $\epsilon$, then both $S_{\rm b}$ and $G_{\rm N}$ increase, so the divergence constraint becomes looser while the norm bound becomes tighter. The flux of $v$ through $A$ does not change, but what happens to that flux in the bulk changes, as less of it can go through the QES and more of it must jump across.
@@ 2.7 Theorems & proofs
In this subsection, we state and prove several theorems that were used in the rest of the section. These proofs use the same tools of convex analysis, especially convex relaxation and dualization, as in the Riemannian max flow-min cut theorem and related theorems, which may be found in Headrick and Hubeny (2018). Therefore, we will be brief. We will also not make any attempt to be rigorous, particularly as regards the measure theory and functional analysis aspects of the proofs.
@@ 2.7.2 Theorems
Lagrange duality will be used in many of our proofs to establish equivalence of optimisation problems. To aid the reader, let us give a one paragraph summary (see the textbook Boyd and Vandenberghe (2004) for further details). We start from a optimisation problem called the primal,
$\min_{x}f_{0}(x)\text{ subject to }f_{i}(x)\leq 0\text{ for }(i=1,\dots,m),\,h_{j}(x)=0\text{ for }(j=1,\dots,n),$ (51)
construct a Lagrangian with Lagrange multiplier terms to enforce the primal’s constraints, then optimise over the variables in the primal problem to arrive at the Lagrange dual problem. A sufficient condition for the primal and dual problems to be equivalent, i.e. to have to same optimal value, for strong duality to hold, is for the primal problem to be convex and for Slater’s condition to be satisfied. This condition requires there to exist a strictly feasible point in the domain of the primal problem, meaning a point that strictly satisfies the nonlinear constraints.
@@
The proof is similar to that in Rolph (2023) for the corresponding theorem for loose flows. We write the RHS of (68) as a concave program:
@@
This theorem was proved in Rolph (2023). Here, we give an alternate proof that involves dualizing from surfaces to flows, similar to the second proof of theorem 2.4. On a technical note, this proof requires an additional assumption, that there is a large volume near $\partial\Sigma$, such as when $\Sigma$ is the region within a large radial cutoff surface of an asymptotically hyperbolic manifold.
@@ 2.8 Multiflows
From the definition, each component vector field $v_{ij}$ is itself a classical flow. More than that, any linear combination $\sum_{i<j}c_{ij}v_{ij}$ where the coefficients obey $|c_{ij}|\leq 1$ is a classical flow. In particular, for any individual boundary region $A_{i}$, we can define the flow $v_{i}:=\sum_{j}v_{ij}$; its flux $\int_{A_{i}}v_{i}$ is bounded above by $S(A_{i})$. The following theorem, proved in Cui et al. (2019), shows that there exists a classical multiflow that simultaneously saturates these bounds for all $i$:
@@ 3.2.2 The strict constraint is saturated on the entanglement wedge
We can actually show something more general that implies what we want: if $\int_{A}v=S_{\rm gen}({r})$, for any $\eth r$} homologous to $A$, then (113) is saturated for that $\eth r$}, and $v$ is perpendicular to $\eth r$}. To show this, we apply (113) to any $\eth r$} homologous to $A$ and with $\int_{A}v=S_{\rm gen}({r})$ gives
$\int_{\eth r}|v|+\left|\int_{\eth r}v-S_{\rm gen}({r})\right|\leq S_{\rm gen}({r}).$ (118)
@@
$\int_{\eth r}|v|+\int_{\eth r}v\leq 2S_{\rm gen}({r}).$ (120)
If either one of these is saturated then (118) must be saturated. The inequality (119) together with the identity $\int_{\eth r}|v|-\int_{\eth r}v\geq 0$ imply that (119) is saturated, and therefore that (1) $v$ on $\eth r$} is parallel to the normal, and (2) that (118) is saturated.
@@ 3.2.3 Constraints implied by the strict constraint
$\begin{split}|v(x)|&\leq\inf_{\eth r\ni x}\frac{S_{\rm gen}(r)}{|\eth r|}\\ &=\frac{1}{4G_{N}}+DS_{b}(x),\qquad DS_{b}(x)=\inf_{\eth r\ni x}\frac{S_{b}(r)}{|\eth r|}\end{split}$ (125)
@@ 3.3 Divergenceless, loose, cutoff-independent flows
The divergenceless, loose, cutoff-independent prescription is (111). These divergenceless flows have similarities and differences with respect to the strict flows. The divergenceless flow lines cannot start or end at points in the bulk. From their definition, the flows obey (123), just like the strict flows, but only for $r\in\mathcal{R}_{A}$. The divergenceless flows do obey the same local norm bound (125) as the strict flow, but we will have to derive it another way, because in this prescription we cannot apply (123) to arbitrary bulk regions.
@@ 3.4.1 Loose flow prescription
$\sup_{v}\int_{A}v\qquad\text{subject to }\int_{{\eth r}}|v|-\int_{r}\nabla\cdot v\leq S_{\rm gen}({r})$ (129)
@@
$L[v,\mu]=\int_{A}v-\int_{\partial\Sigma}\psi v-\int_{\mathcal{R}_{A}}d\mu\left(\int_{\Sigma}\left(d\chi_{r}\cdot v+|v||d\chi_{r}|\right)-S_{\rm gen}(r)\right),\qquad\psi:=\int_{\mathcal{R}_{A}}d\mu\chi_{r}.$ (131)
@@ 3.4.3 Strict flow prescription: surface to flow dualisation
Here we will dualise from the QES prescription to the strict flow prescription. Since we already proved the equivalence between the prescriptions in section 3.4.2, this derivation is redundant, but we include it to show how it can be done because surface-to-flow dualisations are needed when deriving covariant prescriptions Headrick et al. .
@@ 3.4.4 Divergenceless loose flow prescription
If $A$ has a boundary, then (148) is equivalent to the QES prescription. To show this, it is useful to note that an implication of the constraints in (148) is
@@
If $A$ has a boundary, then $|d\chi_{A}|$ is a delta-function on that boundary. Then, from (149), we must have $\int_{\mathcal{R}_{A}}d\mu\geq 1$, and so (148) is lower bounded by $S(A)$. But it is also upper bounded by $S(A)$, because $\lambda=\chi_{r_{QES}}$ and $\mu(r)=\delta(r-r_{QES})$ is always feasible. So, (148) gives $S(A)$, if $A$ has a boundary.
@@ 4 Quantum thread distributions
Thread distributions, introduced in Headrick and Hubeny (2023), are an alternative mathematical representation of bit threads, replacing the flow $v$ with a measure $\upmu$ over bulk curves connecting boundary points. Although they were defined both in the static and covariant settings, here we will focus (as we do throughout this paper) with the static case, and explain a possible method to reproduce the QES formula using thread distributions. Unfortunately, we will have to leave several key statements as conjectures.
@@ 4.1 Entanglement pair functions
In upgrading classical TDs to quantum ones, we will need the concept of an entanglement pair function, which represents the bulk entropies and plays the same role for quantum TDs that the entanglement density function (EDF), defined and studied in subsection 2.3 and section 5, plays for quantum flows. We define an entanglement pair function (EPF) as a non-negative function $g$ on $\Sigma^{2}:=\Sigma\times\Sigma$ such that [^12] Because our application of EPFs will be in holography, in defining EPFs, we are using the notation corresponding to bulk entropies in holography. However, the concept can be applied to any state of a field theory for which entropies of spatial regions are defined, or even more generally, to any multiparty quantum state. Note also that (152) is a property shared by the so-called two-point partial entanglement entropy $\mathcal{I}(x,y)$, defined in Lin et al. (2024).
@@ 4.4 Double holography
In the double holography setting, making use of the above mapping from classical TDs to EPFs, we can also prove conjecture 4.1. Namely, starting from an EDF $f$ on $\Sigma$, we lift it to a classical flow $\tilde{v}$ in $\tilde{\Sigma}$ using theorem 2.2; we then map $\tilde{v}$ to a classical TD $\tilde{\upmu}$ on $\tilde{\Sigma}$ via the flow lines of $\tilde{v}$, as described in subsection 5.1.1 of Headrick and Hubeny (2023) (except that, rather than restricting to threads connecting specified boundary regions, we keep all threads connecting any two boundary points, and we retain their orientations); finally, using (173) we map $\tilde{\upmu}$ to an EPF $g$. Through this series of transformations, $f(x)$ becomes the boundary flux $\tilde{n}\cdot\tilde{v}(x)$, which in turn becomes the number of threads beginning (if $f(x)>0$) or ending (if $f(x)<0$) at $x$, finally implying that $g$ obeys $\int_{\Sigma}dy\,g(x,y)=|f(x)|$.
@@ 5 The entropohedron
In the definition of strict quantum flows given in section 2, the concept of entanglement distribution function (EDF) played a key role. In that setting, the relevant entropies were those of spatial regions of the bulk in a holographic spacetime. The notion of EDF is also applicable for a quantum state on a finite number of parties. For $N$ parties, the EDFs elegantly repackage the entropy function, which is a function of $2^{N}-1$ variables (or equivalently the entropy vector, a vector in $2^{N}-1$ dimensions), into a convex polytope in just $N$ dimensions, which we will call the entropohedron. Whereas the entropy function treats all sets of parties on an equal footing, the entropohedron makes use of the local structure, i.e. which sets of parties are contained within which other sets. It also naturally builds in important properties of entropies such as strong subadditivity. In this section, we give a self-contained presentation of the entropohedron and its properties. In this finite-dimensional setting, we will be able to prove the relevant properties rigorously, as no analysis is required.
Some aspects of this discussion appeared previously in the original bit thread paper Freedman and Headrick (2017). The entropohedron is also closely related to several objects that have been previously studied in the context of submodular functions, including the so-called symmetric submodular polytope (see Bach and others (2013)) and the generalized polymatroid (see Frank et al. (2014)). The simple relation between these objects and network flows, discussed in subsection 5.6, has not been noted before as far as we know. The entropohedron is also related to the notion of entanglement contour, as we will discuss in subsection 5.3.
@@
5.1 Definitions & examples
Everywhere in this section, $X$ will denote a finite set and $2^{X}$ its power set (the set of all subsets of X{X}). We will use $x,y,\ldots$ to represent elements of X{X} and $A,B,\ldots$ to represent subsets of $X$. We denote by $\mathbf{R}^{X}$ the |X||{X}|-dimensional real vector space with basis vectors labelled by elements of X{X}.
@@ 5.1 Definitions & examples
An entropy function [^16] SS is also sometimes called an entropy vector, as it may be considered a vector in $\mathbf{R}^{2^{X}\setminus\{\emptyset\}}$, where we remove the empty set from the basis since we require $S(\emptyset)=0$., [^17] The strong subadditivity condition (180) is also called submodularity. In the literature on submodular functions, in particular in defining the symmetric submodular polytope, the condition (179) is also usually imposed, along with monotonicity, $S(AB)\geq S(A)$ Bach and others (2013). Monotonicity is obeyed by the entropies of classical probability distributions; however, for quantum states it is replaced by the weak monotonicity condition (181). for X{X} is a function $S:2^{X}\to[0,\infty)$, satisfying the following conditions: [^18] Note that (181), applied to the case $B=A$, together with (179), implies $S(A)\geq 0$.
@@
$\forall\,A,B\subseteq X\,,\quad S(A)+S(B)$ $\geq S(A\cap B)+S(A\cup B)\qquad\text{(SSA)}$ (180)
$\forall\,A,B\subseteq X\,,\quad S(A)+S(B)$ $\geq S(A\setminus B)+S(B\setminus A)\qquad\text{(WM)}\,.$ (181)
The entropies of a multipartite quantum system in a fixed state obey (179)–(180), hence the name “entropy function”. [^19] The name “entropy function” is a bit of an abuse, since not every entropy function represents the entropies of a quantum state. In particular, it is known that the entropies of quantum states on at least four parties obey a further set of constrained inequalities Linden and Winter (2005); Cadney et al. (2012). For this reason, we will sometimes use the language of “states” and “parties”. The only aspects of the state that are relevant for our purposes, however, are the entropies. Another natural source of functions obeying these properties is minimal cuts on weighted graphs. We will discuss these further in subsection 5.6. Of course, minimal cuts on graphs are related to quantum states via the RT formula.
@@
Given an entropy function $S$ for $X$, an entanglement distribution function (EDF) is a function $f:X\to\mathbf{R}$ such that, for all $A\subseteq X$,
@@
Given an entropy function $S$, the entropohedron $F_{S}$ is the set of EDFs of $S$.
Figure 11: Entropohedra for various states on three parties: (a) maximally mixed. (b) Maximally classically correlated. (c) Maximally entangled (e.g. GHZ or W state). (d) Marginal of four-party perfect tensor. Axes are $f(1),f(2),f(3)$. Figures are scaled to have equal single-party entropies. [^21] To manipulate these 3D figures, paste the following commands into Mathematica:
RegionPlot3D[Abs[f1] ¡= 1 Abs[f2] ¡= 1 Abs[f3] ¡= 1,—
f1, -1.1, 1.1, f2, -1.1, 1.1, f3, -1.1, 1.1, PlotPoints -¿ 100]—
RegionPlot3D[Abs[f1] ¡= 1 Abs[f2] ¡= 1 Abs[f3] ¡= 1 —
Abs[f1 + f2] ¡= 1 Abs[f2 + f3] ¡= 1 Abs[f1 + f3] ¡= 1 Abs[f1 + f2 + f3] ¡= 1, —
f1, -1.1, 1.1, f2, -1.1, 1.1, f3, -1.1, 1.1, PlotPoints -¿ 100]—
RegionPlot3D[Abs[f1] ¡= 1 Abs[f2] ¡= 1 Abs[f3] ¡= 1 —
Abs[f1 + f2] ¡= 1 Abs[f2 + f3] ¡= 1 Abs[f1 + f3] ¡= 1 Abs[f1 + f2 + f3] ¡= 0.05,—
f1, -1.1, 1.1, f2, -1.1, 1.1, f3, -1.1, 1.1, PlotPoints -¿ 100]—
RegionPlot3D[Abs[f1] ¡= 1 Abs[f2] ¡= 1 Abs[f3] ¡= 1 Abs[f1 + f2 + f3] ¡= 1,—
f1, -1.1, 1.1, f2, -1.1, 1.1, f3, -1.1, 1.1, PlotPoints -¿ 100]—
Examples of entropohedra for various quantum states on two and three parties are shown in Figs. 10 and 21 respectively.
5.2 Basic properties
Proofs of the propositions in this subsection are straightforward, and we leave them to the reader. The following properties of the entropohedron follow directly from the definition: